2. (i) Verify that
u(x,y;λ) = e−λy cosλx, −∞ < λ < ∞
is a one parameter family of solutions of Laplaces’s equation in R2.
(ii) Find v(x,y;λ) = ∂ ∂λu(x,y;λ) and verify that v(x,y;λ), − ∞ < λ < ∞, is also a one parameter family of solutions of Laplace’s equation in R2.
(iii) For (x,y) in the upper half plane y > 0, the improper integral
v(x,y) = ∞ u(x,y;λ)dλ 0
is convergent. Evaluate this integral and show by direct computation that v(x,y) is a solution of Laplace’s equation in the upper half plane.
1
Expert's answer
2020-09-07T16:46:08-0400

(i) We compute the derivatives and find that

"u_{xx}=-\lambda^2e^{-\lambda y}cos(\lambda x),\quad u_{yy}=\lambda^2e^{-\lambda y}cos(\lambda x)"

From the latter it follows that the Laplace equation is satisfied. I.e., "u_{xx}+u_{yy}=0" .

(ii)

"u_{\lambda}=-\lambda e^{-\lambda y}cos(\lambda x)-e^{-\lambda y}sin(\lambda x),"

"u_{\lambda\lambda}=\lambda^2e^{-\lambda y}cos(\lambda x)+\lambda^2e^{-\lambda y}sin(\lambda x)+\lambda e^{-\lambda y}sin(\lambda x)-\lambda e^{-\lambda y}cos(\lambda x)="

"=\lambda^2u-\lambda u+\lambda^2 v+\lambda v,"

where "v=e^{-\lambda y}sin(\lambda x)" . We compute the second order derivatives and get "v_{xx}=-\lambda^2e^{-\lambda y}sin(\lambda x), v_{yy}=\lambda^2e^{-\lambda y}sin(\lambda x)" .

Thus, "u_{\lambda\lambda}" also satisfies the Laplace equation.

(iii) At first, we calculate the indefinite integral "\int e^{-\lambda y}cos(\lambda x)d\lambda=-\frac{1}{y}e^{-\lambda y}cos(\lambda x)+"

"-\frac{x}{y}\int e^{-\lambda y}sin(\lambda x)d\lambda=-\frac{1}{y}e^{-\lambda y}cos(\lambda x)+\frac{x}{y^2}e^{-\lambda y}sin(\lambda x)-\frac{x^2}{y^2}\int e^{-\lambda y}cos(\lambda x)d\lambda"

We denote "I=\int e^{-\lambda y}cos(\lambda x)d\lambda" . Then "I=-\frac{1}{y}e^{-\lambda y}cos(\lambda x)+\frac{x}{y^2}e^{-\lambda y}sin(\lambda x)-\frac{x^2}{y^2}I" .

From the latter we get: "I=-\frac{y}{x^2+y2}e^{-\lambda y}cos(\lambda x)+\frac{x}{x^2+y^2}e^{-\lambda y}sin(\lambda x)"

Using the latter, we get "v(x,y)=\int_0^{+\infty} e^{-\lambda y}cos(\lambda x)d\lambda=\frac{y}{x^2+y^2}"

The latter has the following derivatives:

"v_{xx}=\frac{2y(3x^2-y^2)}{(y^2+x^2)^3};\qquad v_{yy}=-\frac{2y(3x^2-y^2)}{(y^2+x^2)^3}"

From the latter it is clear that "v_{xx}+v_{yy}=0" .